OpenStudy (anonymous):

the area of a triange can be modeled: A(x)=x^3-1. The length is x-1. Find a polynomial to represent the width of the triangle. (the answer is x^2 + x +1 .... how do u get that?)

OpenStudy (anonymous):

Is this an equilateral triangle? or an arbitrary triangle?

OpenStudy (watchmath):

Notice that \(x^3-1=(x-1)(x^2+x+1)\) So the width od the triangel is \(2(x^2+x+1)\)

OpenStudy (anonymous):

not sure why the "2" is there. \[\frac{x^3-1}{x-1}=x^2+x+1\]

OpenStudy (watchmath):

I guess area of triangle is 1/2*base*height

OpenStudy (anonymous):

is the length and width? can we base and sides?

OpenStudy (anonymous):

oops you are right. how could i doubt it?

OpenStudy (anonymous):

OMG I MENT RECTANGLE SRY GUYS

OpenStudy (anonymous):

lol well i read "rectangle' too so it must have been a subliminal message. then there is no 2!

OpenStudy (anonymous):

x^3−1 -------- x−1 =x^2+x+1 how is that equal to that?

OpenStudy (anonymous):

actually, you are just going to use the formula of the area of a rectangle which is length times width A=lw \[x ^{3}-1=w(x-1)\] divide both sides by x-1 then you will get x^3−1 -------- x−1 which is equal to x^2+x+1 through factoring its gcf.